# Vulnerabilities and capabilities in the EU Automotive industry: Leveraging Input-Output Analysis and Economic Complexity

### v1

@@ Abstract
Keywords: automotive, input-output, global value chains, economic fitness and complexity, relatedness
@@ 1 Introduction
These trends are increasingly straining the EU automotive industry, prompting policymakers to propose provisional tariffs on electric vehicles imported from China. [^3] https://ec.europa.eu/commission/presscorner/detail/en/statement_24_5041 This initiative aligns with the broader framework of Open Strategic Autonomy, developed after the COVID-19 pandemic in response to supply chain disruptions and shortages. The framework aims at achieving a new balance between security and competitiveness – preserving trade openness while mitigating vulnerabilities from economic interdependence. This ”de-risking” approach seeks to mitigate vulnerabilities and supply chain interdependences while maintaining an open global economy (Van den Abeele,, 2021; Gehrke,, 2022; Schmitz and Seidl,, 2023).
Against the backdrop of supply chain contractions and the electric mobility transition, this paper employs input-output analysis, Economic Complexity methods, and machine learning to investigate the EU27 automotive industry’s input dependencies, competitiveness, and vulnerabilities.
To fully grasp the transformations reshaping the sector, the study moves beyond traditional sectoral perspectives by accounting for the global fragmentation of production along value chains. The automotive industry relies on the assembly of thousands of components sourced globally, exposing it to strategic dependencies and supply chain vulnerabilities. A particularly significant example is electric vehicle batteries, the majority of which are currently imported from China.
To address these challenges, we propose a novel empirical strategy that simultaneously analyzes supply chain dependencies and export capabilities. This approach integrates product-level trade data with inter-sectoral linkages—two dimensions rarely examined together in supply chain studies. Additionally, the methodology incorporates a microdata-validated list of 63 harmonized system (HS) intermediate products related to motor vehicle production, provided in the Appendix, enabling granular and precise analysis.
@@
To fully capture the transformations reshaping the sector, the study moves beyond conventional sectoral perspectives by recognizing that the EU automotive industry’s productive capabilities are shaped by the global fragmentation of production along value chains. Manufacturing a car involves assembling thousands of components sourced worldwide, exposing the industry to strategic dependencies and supply chain vulnerabilities. As mentioned above, a case in point is electric vehicle batteries, a critical component predominantly sourced through international supply chains, with a significant share currently imported from China. To address these challenges, we propose a novel empirical strategy that simultaneously analyzes supply chain dependencies and export capabilities. This approach integrates product-level trade data with inter-sectoral linkages—two dimensions rarely examined together in supply chain studies. Additionally, the methodology incorporates a microdata-validated list of 63 harmonized system (HS) intermediate products related to motor vehicle production, provided in the Appendix, enabling granular and precise analysis.
@@
In our empirical analysis, first, we examine the industry’s dependence on key trade partners and products, highlighting a shift in suppliers through input-output analysis and product-level trade . China, along with Turkey, South Korea, and Mexico, has gained importance, while countries like the UK and Japan have seen a decline in relevance. Notably, imports of electrical parts, especially electric accumulators, have surged, with China as the EU’s primary supplier. Second, using the Economic Fitness and Complexity metric, we evaluate the competitiveness of EU27 countries in automotive-related exports. Our analysis shows that Eastern European nations—particularly Poland, Hungary, and Czechia—have improved their competitiveness in automotive components, surpassing Germany. Additionally, these countries, along with Italy, display the highest probability of developing future comparative advantages in the automotive components. Third, we focus on the electric mobility transition: while EU27 exports of electric and hybrid vehicles have grown, internal combustion engine vehicles still dominate, and the region lags behind in battery production capabilities, remaining heavily reliant on lithium-ion battery imports from China. Some EU countries are positioned to become competitive in certain types of batteries, though very few are positioned to become competitive in lithium-ion battery exporters by 2027. Finally, we conduct a vulnerability assessment of the EU27 automotive supply chain, combining product-level data on net exposure to extra-EU27 inputs and supply chain import concentration. Products in the Electrical and Electric Parts and Engines and Parts categories are identified as the most vulnerable, with lithium-ion batteries and electric accumulators posing significant risks due to high external dependency and supply concentration.

### v2

@@ Abstract
Keywords: European automotive industry, electric vehicles, value chain vulnerability, input-output, global value chains, economic fitness and complexity, relatedness
@@ 1 Introduction
These trends are increasingly straining the EU automotive industry, prompting policymakers to propose provisional tariffs on electric vehicles imported from China. [^3] https://ec.europa.eu/commission/presscorner/detail/en/statement_24_5041 This initiative aligns with the broader framework of Open Strategic Autonomy, developed after the COVID-19 pandemic in response to supply chain disruptions and shortages. The framework aims at achieving a new balance between security and competitiveness – preserving trade openness while mitigating vulnerabilities from economic interdependence. A ”de-risking” approach that seeks to mitigate vulnerabilities and supply chain interdependencies while maintaining an open global economy (Van den Abeele,, 2021; Gehrke,, 2022; Schmitz and Seidl,, 2023).
@@
Against the backdrop of supply chain contractions and the electric mobility transition, this paper employs input-output analysis, Economic Complexity, and machine learning to investigate the EU27 automotive industry’s input dependencies, competitiveness, and vulnerabilities. To fully capture the sector’s transformations, the study moves beyond conventional sectoral perspectives by recognizing that the EU automotive industry’s productive capabilities are shaped by the global fragmentation of production along value chains. Manufacturing a car involves assembling thousands of components sourced worldwide, exposing the industry to strategic dependencies and supply chain vulnerabilities. A case in point is electric vehicle batteries, a critical component predominantly sourced through international supply chains, with a significant share currently imported from China. We thus propose a novel empirical strategy that simultaneously analyses supply chain dependencies and export capabilities integrating granular product-level trade data with inter-sectoral linkages—two dimensions rarely examined together in supply chain studies. To this aim, we propose a novel a microdata-validated list of 63 harmonized system (HS) intermediate products related to motor vehicle production, provided in the Appendix.
@@
In our empirical analysis, first, we examine the industry’s dependence on key trade partners and products, highlighting a shift in suppliers through input-output analysis and product-level trade. China, along with Turkey, South Korea, and Mexico, has gained importance, while countries like the UK and Japan have seen a decline in relevance. Notably, imports of electrical parts, especially electric accumulators, have surged, with China as the EU’s primary supplier. Second, using the Economic Fitness and Complexity metric, we evaluate the competitiveness of EU27 countries in automotive-related exports. Our analysis shows that Eastern European nations—particularly Poland, Hungary, and Czechia—have improved their competitiveness in automotive components, surpassing Germany. Additionally, these countries, along with Italy, display the highest probability of developing future comparative advantages in automotive components. Third, we focus on the electric mobility transition: while EU27 exports of electric and hybrid vehicles have grown, internal combustion engine vehicles still dominate, and the region lags behind in battery production capabilities, remaining heavily reliant on lithium-ion battery imports from China. Some EU countries are positioned to become competitive in certain types of batteries, though very few are positioned to become competitive in lithium-ion battery exporters by 2027. Finally, we conduct a vulnerability assessment of the EU27 automotive supply chain, combining product-level data on net exposure to extra-EU27 inputs and supply chain import concentration. Products in the Electrical and Electric Parts and Engines and Parts categories are identified as the most vulnerable, with lithium-ion batteries and electric accumulators posing significant risks due to high external dependency and supply concentration.